OpenStudy (anonymous):

Determine the second and third therms of an arithmetic sequence if: The First term is 42 and the fourth term is 27

OpenStudy (whpalmer4):

an arithmetic sequence is one where successive terms differ by a fixed amount (the common difference). You have the first term = 42, and the fourth term = 27. You need to find the second and third terms. 42, x, y, 27 42-x = x-y = y-27 or if a is the common difference, 42 - a = x 42 - 2a = y 42 - 3a = 27 you should be able to find a from the last equation, and then figure out the middle two terms.

OpenStudy (anonymous):

I still don't get it.... like at all

OpenStudy (anonymous):

i have to use the formula tn=t1+(n-1)d

OpenStudy (whpalmer4):

okay, fine. Your formula is the equivalent of mine. \[t_n = t_1+(n-1)d\]\(t_n\) is the \(n\)th term \(t_1\) is the 1st term \(n\) is the term you seek \(d\) is the common difference (which you don't know yet) We know \(t_1=42 \text{ and }t_4 = 27\)Can you fill those into the equation and tell me what you get after doing so?

OpenStudy (whpalmer4):

use \(n=4\) because you are using \(t_4\)

OpenStudy (anonymous):

ok--- \[42= 27+ (4-1)d\] Is that right?

OpenStudy (whpalmer4):

Isn't the first term 42, and the 4th term 27?

OpenStudy (whpalmer4):

you have them switched.

OpenStudy (anonymous):

Oh woops sorry let me try again

OpenStudy (anonymous):

\[27= 42+(4-1)d\]

OpenStudy (whpalmer4):

Good, can you solve that for \(d\)?

OpenStudy (anonymous):

so then d= 45?

OpenStudy (whpalmer4):

show me your work that led to that answer

OpenStudy (anonymous):

I just did 42+3 then that would equal 45 right?

OpenStudy (whpalmer4):

what happened to \(d\)? and the 27 on the left?

OpenStudy (anonymous):

I'm so confused

OpenStudy (whpalmer4):

you have an equation: \[27 = 42 + (4-1)d\] Read in words, you could see that as "some number I don't know, which I'll call d, times the quantity 4-1, plus an additional 42, equals 27."

OpenStudy (anonymous):

so 27= 45(d)

OpenStudy (whpalmer4):

no! the right side of that equation is \[42 + (4-1)*d\] You can't just add 42 + (4-1) = 45!

OpenStudy (whpalmer4):

What can we do with the \[(4-1)d\]? We can distribute it: \[(4-1)d = 4*d-1*d\]or we can simplify it:\[(4-1)d = (3)d = 3d\] We could also distribute it, then simplify it: \[(4-1)d=4*d-1*d=4d-d=3d\]

OpenStudy (anonymous):

Welcome to OpenStudy @Animegurl3 !

OpenStudy (anonymous):

OH!! I get it now... well the (4-1)d part so just simplifying makes the 3 the difference right?

OpenStudy (whpalmer4):

uh, no. the value of \(d\) is the difference. We still need to find that. \[27=42+(4-1)d\]we need to manipulate this equation so that we end up with just \(d\) on one side of the equals sign and a number on the other. You can add, subtract, multiply, divide, etc so long as you do the same thing to both sides of the equals sign.

OpenStudy (anonymous):

oh okay so then 27=42+(4-1)d which becomes \[27=3d\] and then you divide 3 from both sided so then 9=d right?

OpenStudy (anonymous):

Hey are you still there?

OpenStudy (whpalmer4):

No, I'm banging my head on the desk :-) \[27 = 42+(4-1)d\]How did you get from there to \[27=3d\]What happened to the \( 42\)?!?

OpenStudy (anonymous):

Oh funny :). Sorry! Umm... then 27=42+3d then subtract 42 from both sides and get -15=3d then divide 3 from both sides and -5=d HA That HAS to be right

OpenStudy (whpalmer4):

yes, much better. Now what happens if you add -5 to 42? What do you get? And if you add -5 to that number, what do you get? And if you add -5 to that number, what do you get? Discuss.

OpenStudy (anonymous):

okay so then t1= 42 and t2=37 and t3=32 and t4=27 OMG I get Thank You soo MUCH

OpenStudy (anonymous):

I GET IT I GET IT Thanx for having soo much patience with me :-) :-)

OpenStudy (anonymous):

Dunno how he does it.

OpenStudy (anonymous):

If only my teacher made this much sense. She basically mutters through the whole thing and gets angry when nobody gets it like in 3 seconds

OpenStudy (whpalmer4):

it all balances out, I have 0 patience with the people who actually have to live with me :-)

OpenStudy (anonymous):

haha :)

OpenStudy (whpalmer4):

So, what is the 7th term of this sequence?